# The Covariant Stone-von Neumann Theorem for Locally Compact Quantum Groups

### v1

@@ 4. The Covariant Stone-von Neumann Theorem
Note that since $A$ is separable, it has a faithful state. Let $\vartheta_{2},\tilde{\vartheta_{2}}$ be irreducible representations of $A$ and $\vartheta_{1}$ an irreducible representation of $\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes C_{0}^{u}(\widehat{\mathbb{G}})$. Then we can form a representation $\vartheta_{1}\otimes\vartheta_{2}$ of $(\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes C_{0}^{u}(\widehat{\mathbb{G}}))\otimes_{\textnormal{max}}A$ on $H_{\vartheta_{1}}\otimes H_{\vartheta_{2}}$; same for $\vartheta_{1}\otimes\tilde{\vartheta}_{2}$. They are irreducible, hence by isomorphism (7) and Proposition 4.5, there is a unitary $v\colon H_{\vartheta_{1}}\otimes H_{\vartheta_{2}}\rightarrow H_{\vartheta_{1}}\otimes H_{\tilde{\vartheta}_{2}}$ such that $v(\vartheta_{1}\otimes\vartheta_{2})(\cdot)=(\vartheta_{1}\otimes\tilde{\vartheta}_{2})(\cdot)v$. Extend these representations to multiplier algebras, and consider representations $A\ni a\mapsto(\vartheta_{1}\otimes\vartheta_{2})(1\otimes a)=1\otimes\vartheta_{2}(a)\in\mathbb{B}(H_{\vartheta_{1}}\otimes H_{\vartheta_{2}})$, and same for $\tilde{\vartheta}_{2}$. These are multiples of $\vartheta_{2},\tilde{\vartheta}_{2}$ which are equivalent via the unitary $v$. Since $\vartheta_{2},\tilde{\vartheta}_{2}$ are irreducible, we can use again [DixmierC, Proposition 5.3.3] to deduce that $\vartheta_{2},\tilde{\vartheta}_{2}$ are equivalent. We assume that $A$ is separable, hence [DixmierC, Exercise 4.7.2] tells us that $A$ is elementary.
@@
Let $(\mathbb{G},A,\alpha)$ be a left coaction, where $\mathbb{G}$ is a regular locally compact quantum group and $A$ a $C^{*}$-algebra which admits a faithful state. If $(\mathbb{G},A,\alpha)$ is subject to Stone-von Neumann uniqueness, then $\alpha$ is a maximal coaction.
@@
We have a canonical surjective ***-homomorphism $\rho\colon\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A)\rightarrow\mathbb{K}(L^{2}(\mathbb{G}))\otimes A$, and we want to show that this map is injective. Assume otherwise, i.e. assume that $\text{ker}(\rho)$ is a nontrivial ideal in $\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A)$. Then we can form quotient $C^{*}$-algebra $(\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A))/\text{ker}(\rho)$ and choose its irreducible representation $\theta\colon(\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A))/\text{ker}(\rho)\rightarrow\mathbb{B}(H_{\theta})$. Composing with the quotient map, we obtain an irreducible representation $\tilde{\theta}\colon\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A)\rightarrow\mathbb{B}(H_{\theta})$ such that $\text{ker}(\rho)\subseteq\text{ker}(\tilde{\theta})$. On the other hand, since $\text{ker}(\rho)$ is nontrivial, we can choose $x\in\text{ker}(\rho)\setminus\{0\}$ and irreducible representation $\tilde{\theta}^{\prime}\colon\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A)\rightarrow\mathbb{B}(H_{\tilde{\theta}^{\prime}})$ such that $\tilde{\theta}^{\prime}(x)\neq 0$. This shows that $\tilde{\theta},\tilde{\theta}^{\prime}$ have different kernels, in particular are not equivalent – this gives us a contradiction with Proposition 4.5. Thus $\rho$ is an isomorphism, and the coaction $(\mathbb{G},A,\alpha)$ is maximal.
As a final application, we identify another sufficient condition for strong regularity of quantum group $\widehat{\mathbb{G}}$.
@@
Let $\mathbb{G}$ be a regular locally compact quantum group and assume that $(\mathbb{G},\mathbb{C},\operatorname{triv})$ is Stone-von Neumann unique. Then $\widehat{\mathbb{G}}$ is strongly regular.
@@
The previous corollary gives us an isomorphism $\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes C_{0}^{u}(\widehat{\mathbb{G}})\rightarrow\mathbb{K}(L^{2}(\mathbb{G}))$. Let $j_{{\widehat{\mathbb{G}}}^{\textnormal{op}}}\colon C_{0}^{u}(\mathbb{G}^{c})\rightarrow M({\widehat{\mathbb{G}}}^{\textnormal{op}}\ltimes C_{0}(\widehat{\mathbb{G}}))$ be the canonical inclusion. Observe that $(j_{{\widehat{\mathbb{G}}}^{\textnormal{op}}},j_{C_{0}(\widehat{\mathbb{G}})}\hat{\lambda})$ is a covariant representation of the system $(\widehat{\mathbb{G}}^{\textnormal{op}},C_{0}^{u}(\widehat{\mathbb{G}}),(\hat{\lambda}\otimes\text{id})\circ\Delta^{u}_{\widehat{\mathbb{G}}^{\textnormal{op}}})$ on $\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes C_{0}(\widehat{\mathbb{G}})$ whose integrated form determines a (nonzero) surjection of $\mathbb{K}(L^{2}(\mathbb{G}))$. Since $\mathbb{K}(L^{2}(\mathbb{G}))$ is simple, it has to be an isomorphism. We conclude that $\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes C_{0}(\widehat{\mathbb{G}})$ comprises compact operators so that by definition, $\widehat{\mathbb{G}}^{\textnormal{op}}$ is strongly regular. This property is equivalent to strong regularity of $\widehat{\mathbb{G}}$, which ends the proof. ∎

### v2

@@ 4. The Covariant Stone–von Neumann Theorem
Note that since $A$ is separable, it has a faithful state. Let $\vartheta_{2},\tilde{\vartheta_{2}}$ be irreducible representations of $A$ and $\vartheta_{1}$ an irreducible representation of $\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes C_{0}^{u}(\widehat{\mathbb{G}})$. Then we can form a representation $\vartheta_{1}\otimes\vartheta_{2}$ of $(\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes C_{0}^{u}(\widehat{\mathbb{G}}))\otimes_{\textnormal{max}}A$ on $H_{\vartheta_{1}}\otimes H_{\vartheta_{2}}$; same for $\vartheta_{1}\otimes\tilde{\vartheta}_{2}$. They are irreducible, hence by the isomorphism (10) above and Theorem 4.6, there is a unitary $v\colon H_{\vartheta_{1}}\otimes H_{\vartheta_{2}}\rightarrow H_{\vartheta_{1}}\otimes H_{\tilde{\vartheta}_{2}}$ such that $v(\vartheta_{1}\otimes\vartheta_{2})(\cdot)=(\vartheta_{1}\otimes\tilde{\vartheta}_{2})(\cdot)v$. Extend these representations to multiplier algebras, and consider representations $A\ni a\mapsto(\vartheta_{1}\otimes\vartheta_{2})(1\otimes a)=1\otimes\vartheta_{2}(a)\in\mathbb{B}(H_{\vartheta_{1}}\otimes H_{\vartheta_{2}})$, and same for $\tilde{\vartheta}_{2}$. These are multiples of $\vartheta_{2},\tilde{\vartheta}_{2}$ which are equivalent via the unitary $v$. Since $\vartheta_{2},\tilde{\vartheta}_{2}$ are irreducible, we can use again [DixmierC, Proposition 5.3.3] to deduce that $\vartheta_{2},\tilde{\vartheta}_{2}$ are equivalent. We assume that $A$ is separable, hence [DixmierC, Exercise 4.7.2] tells us that $A$ is elementary. ∎
@@
Let $(\mathbb{G},A,\alpha)$ be a left dynamical system, where $\mathbb{G}$ is a regular locally compact quantum group and $A$ a $C^{*}$-algebra which admits a faithful state. If $(\mathbb{G},A,\alpha)$ is subject to Stone–von Neumann uniqueness, then $\alpha$ is a maximal action.
@@
We have a canonical surjective *-homomorphism $\rho\colon\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A)\rightarrow\mathbb{K}(L^{2}(\mathbb{G}))\otimes A$, and we want to show that this map is injective. Assume otherwise, i.e. assume that $\text{ker}(\rho)$ is a nontrivial ideal in $\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A)$. Then we can form the quotient $C^{*}$-algebra $(\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A))/\text{ker}(\rho)$ and choose its irreducible representation $\theta\colon(\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A))/\text{ker}(\rho)\rightarrow\mathbb{B}(H_{\theta})$. Composing with the quotient map, we obtain an irreducible representation $\tilde{\theta}\colon\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A)\rightarrow\mathbb{B}(H_{\theta})$ such that $\text{ker}(\rho)\subseteq\text{ker}(\tilde{\theta})$. On the other hand, since $\text{ker}(\rho)$ is nontrivial, we can choose $x\in\text{ker}(\rho)\setminus\{0\}$ and an irreducible representation $\tilde{\theta}^{\prime}\colon\widehat{\mathbb{G}}^{\textnormal{op}}\ltimes(\mathbb{G}\ltimes A)\rightarrow\mathbb{B}(H_{\tilde{\theta}^{\prime}})$ such that $\tilde{\theta}^{\prime}(x)\neq 0$. This shows that $\tilde{\theta},\tilde{\theta}^{\prime}$ have different kernels, in particular are not equivalent – this gives us a contradiction with Theorem 4.6. Thus $\rho$ is an isomorphism, and the action $\alpha$ is maximal. ∎
Remark.
Corollaries 4.7 and 4.8 indicate that the assumptions of Theorem 4.3, namely elementariness of the coefficient algebra and maximality of the action, are not conditions designed to simplify arguments. On the contrary, they appear necessary.
As a final application, we prove the converse to Corollary 4.4 and show that Stone–von Neumann uniqueness of $(\mathbb{G},\mathbb{C},{\rm triv})$, where $\mathbb{G}$ is regular, implies maximality of the trivial action of $\mathbb{G}$ on $\mathbb{C}$.
@@
Let $\mathbb{G}$ be a regular locally compact quantum group. The following properties are equivalent:
  1. (1)
    $(\mathbb{G},\mathbb{C},\operatorname{triv})$ is Stone–von Neumann unique,
  2. (2)
    the trivial action of $\mathbb{G}$ on $\mathbb{C}$ is maximal,
  3. (3)
    $\mathbb{G}$ is strongly regular.
@@
If $(\mathbb{G},\mathbb{C},{\rm triv})$ is Stone–von Neumann unique, then triv is maximal by Corollary 4.8. The converse implication is the content of Corollary 4.4. Equivalence of $(2)$ and $(3)$ was proved in Proposition 2.2. ∎